# Understanding How Psychological Distance Influences User Preferences in Conversational Versus Web Search

### v1

@@ 4.1.3. Increase in Relative Perceived Usefulness (H3 Supported)
Figure 5. Comparative analysis of perceived usefulness between WebSearch and ConvSearch at varying psychological distances. Panels (a)-(d) show minimal differences at a near psychological distance, except for social distance. Panels (e)-(h) depict significant increases in perceived usefulness for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived usefulness for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived usefulness at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.337), marked as non-significant (ns); (b) Temporal Distance (p = 0.180), marked as non-significant (ns); (c) Social Distance (p = 0.031), marked with one asterisk (*); (d) Hypothetical Distance (p = 0.418), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived usefulness at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.005), marked with two asterisks (**); panels (f) Temporal Distance (p ¡ 0.001), marked with three asterisks (***); (g) Social Distance (p = 0.001), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.003), marked with two asterisks (**).
At a near psychological distance, perceived usefulness was similar in three out of the four dimensions (Fig. 5 (a),(b) and (d)) between WebSearch (Spatial: $M=6.14$, $SD=1.08$; Temporal: $M=5.87$, $SD=1.09$; Hypothetical: $M=5.44$, $SD=1.79$) and ConvSearch (Spatial: $M=5.68$, $SD=1.50$; Temporal: $M=5.49$, $SD=1.59$; Hypothetical: $M=5.90$, $SD=1.48$). These differences were not significant (Spatial: $t(15)=1.07$, $p=0.300$, Cohen’s $d=0.27$; Temporal: $t(15)=1.35$, $p=0.198$, Cohen’s $d=0.34$; Hypothetical: $t(15)=-0.86$, $p=0.405$, Cohen’s $d=-0.21$). However, with social distance, a significantly higher score ($t(15)=-2.38$, $p=0.031$, Cohen’s $d=-0.60$) for ConvSearch ($M=6.04$, $SD=1.04$) was observed compared to WebSearch ($M=5.33$, $SD=1.50$), as shown in Fig. 5 (c).
@@
In contrast, at a far psychological distance, ConvSearch was perceived as significantly more useful in all four dimensions (Fig. 5 (e)-(h)). With spatial distance, ConvSearch ($M=6.14$, $SD=0.57$) was perceived as more useful compared to WebSearch ($M=5.04$, $SD=1.48$), $t(15)=-2.83$, $p=\textbf{0.006}$, Cohen’s $d=-0.71$. Similar phenomena were observed in the temporal ($M=6.63$, $SD=0.64$ vs. $M=5.26$, $SD=1.41$, $t(15)=-3.79$, $p<\textbf{0.001}$, Cohen’s $d=-0.95$), social ($M=6.54$, $SD=0.59$ vs. $M=4.85$, $SD=1.57$, $t(15)=-3.66$, $p=\textbf{0.001}$, Cohen’s $d=-0.92$), and hypothetical ($M=6.46$, $SD=0.90$ vs. $M=4.81$, $SD=1.78$, $t(15)=-3.29$, $p=\textbf{0.002}$, Cohen’s $d=-0.82$) dimensions. Notably, the social distance dimension not only maintained its significance, but also showed an increased effect size, moving from a marginal to a highly significant difference. These findings support H3, demonstrating that as psychological distance increases, so does the perceived usefulness of ConvSearch compared to WebSearch.
4.1.4. Increase in Relative Perceived Enjoyment (H4 Supported)
@@
Figure 6. Comparative analysis of perceived enjoyment between WebSearch and ConvSearch across four psychological distances. Panels (a)-(c) indicate no significant differences at a near psychological distance, while panel (d) shows a marginal preference for ConvSearch. Panels (e)-(h) demonstrate significant increases in perceived enjoyment for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived enjoyment for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived enjoyment at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.450), marked as non-significant (ns); (b) Temporal Distance (p = 0.669), marked as non-significant (ns); (c) Social Distance (p = 0.212), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.018), marked with one asterisk (*). In the second row, panels (e) through (h) illustrate the perceived enjoyment at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.014), marked with one asterisk (*); panels (f) Temporal Distance (p ¡ 0.001), marked with three asterisks (***); (g) Social Distance (p = 0.004), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.001), marked with two asterisks (**).
@@
At a near psychological distance, perceived enjoyment was similar in three out of the four dimensions (Fig. 6 (a)-(c)) between WebSearch (Spatial: $M=5.40$, $SD=1.03$, Temporal: $M=5.13$, $SD=1.28$, Social: $M=5.33$, $SD=1.34$) and ConvSearch (Spatial: $M=5.00$, $SD=1.81$, Temporal: $M=4.98$, $SD=1.78$), Social: $M=5.75$, $SD=1.21$). These differences were not significant (Spatial: $t(15)=0.78$, $p=0.450$, Cohen’s $d=0.19$; Temporal: $t(15)=0.44$, $p=0.669$, Cohen’s $d=0.11$; Social: $t(15)=-1.30$, $p=0.212$, Cohen’s $d=-0.33$). However, in the hypothetical distance dimension there was a significant difference (ConvSearch: ($M=5.67$, $SD=1.22$; WebSearch: ($M=4.71$, $SD=1.50$, ($t(15)=-2.65$, $p=0.018$, Cohen’s $d=-0.66$). suggesting a higher perceived enjoyment of ConvSearch even at a near psychological distance (Fig. 6 (d)).
In contrast, when psychological distance was far, there were significant differences in all four dimensions (Fig. 6 (e)-(h)). ConvSearch produced greater enjoyment than WebSearch (Spatial: $M=5.46$, $SD=1.05$ vs. $M=4.67$,$SD=1.05$,$t(15)=-2.42$, $p=\textbf{0.014}$, Cohen’s $d=-0.60$; Temporal: $M=5.92$, $SD=1.23$ vs. $M=4.73$, $SD=1.41$, $t(15)=-4.73$, $p<\textbf{0.001}$, Cohen’s $d=-1.18$; Social: $M=5.65$, $SD=1.21$ vs. $M=5.65$, $SD=1.36$, $t(15)=-3.03$, $p=\textbf{0.004}$, Cohen’s $d=-0.76$; Hypothetical: $M=6.00$, $SD=1.23$ vs. $M=4.63$, $SD=1.49$, $t(15)=-3.67$, $p=\textbf{0.001}$, Cohen’s $d=-0.92$). Notably, the difference in perceived enjoyment in the hypothetical distance was greater in the far condition compared to the near condition. These results underscore that as psychological distance increases, the perceived enjoyment of using ConvSearch significantly exceeds that of WebSearch across various dimensions, supporting H4.
4.1.5. Increased Preferences for Conversational Search (H5 Supported)
@@
Figure 7. Differential preference for ConvSearch over WebSearch across various psychological distances, shown in two parts: (a) detailed comparison across four dimensions, and (b) aggregated preference at near and far distances. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
The impact of psychological distance on the acceptance of ConvSearch supports H5. When the psychological distance is close (as shown in Fig. 7 (a)), the preference for ConvSearch is already evident, with 75.0% of participants favoring it in the spatial dimension. However, as the distance increases, this preference increases to 87.5%. Similar patterns are observed in other dimensions as psychological distance increases: for temporal distance, the preference increases from 56.2% social distance, from 75.0% to 81.2%; and in hypothetical distance,93.8%. Upon aggregating the data across all dimensions and analyzing the preference for WebSearch and ConvSearch at near and far psychological distances (Fig. 7 (b)), a chi-square test revealed significant differences ( $\chi^{2}(1,N=128)=4.97$, $p=\textbf{0.026}$). Thus, while participants exhibit a preference for ConvSearch even at a near psychological distance, this preference intensifies significantly at a far psychological distance.
@@ 4.2. Thematic Analysis of User Preferences
We also analysed the qualitative responses regarding participants’ preferences for one system over the other. Some of the themes that emerged were unique to a particular system (i.e. only participants who preferred that system mentioned that theme), while others were shared between both systems.
@@ 4.2.1. Unique Themes for Preferring ConvSearch
These themes highlight the distinct advantages of ConvSearch over WebSearch.
Low Workload. Many participants (n = 58) reported that using ConvSearch required less effort compared to WebSearch. They felt that ConvSearch simplified the information gathering process by eliminating the need to visit multiple websites and compile information independently. As one participant (P6) explained: “It was easier to find out what I needed to know and less work on my part to pull it all together.” Thus, reduced workload was one of the reasons for their preference for ConvSearch.
Speed and Efficiency. Participants (n = 43) perceived ConvSearch as faster and more efficient in making travel plans. The efficiency they focused on was whether ConvSearch could quickly provide information while meeting their needs to complete travel plans. As P108 stated, “I would choose the chatbot because it saves me time and is more efficient at making a full-bodied itinerary.
Specific Information Consolidation. Another major reason participants (n = 43) preferred ConvSearch was its ability to quickly consolidate information from various sources for specific queries. P96 said, “The results the chatbot provides is neater compared to the websearch’s results which are just links while the chatbot makes sure to summarize any relevant info and provides it to me.

### v2

@@ 5.1.3. Psychological Distance Increases Differences in Perceived Usefulness (H3 Supported)
Figure 5. Comparative analysis of perceived usefulness between WebSearch and ConvSearch at varying psychological distances. Panels (a)-(d) show minimal differences at a near psychological distance, except for social distance. Panels (e)-(h) depict significant increases in perceived usefulness for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿ 0.05 (ns), p ¡= 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived usefulness for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks from the paired samples statistical tests. In the first row, panels (a) through (d) depict the perceived usefulness at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.386), marked as non-significant (ns); (b) Temporal Distance (p = 0.262), marked as non-significant (ns); (c) Social Distance (p = 0.058), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.366), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived usefulness at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.006), marked with two asterisks (**); panels (f) Temporal Distance (p = 0.004), marked with two asterisks (**); (g) Social Distance (p = 0.005), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.008), marked with two asterisks (**).
At a near psychological distance (as illustrated in Fig. 5 (a)-(c)), the differences in perceived usefulness between the WebSearch (Spatial Distance: $M=6.14$, $SD=1.08$; Temporal Distance: $M=5.87$, $SD=1.09$; Social Distance: $M=5.33$, $SD=1.50$; Hypothetical Distance: $M=5.44$, $SD=1.79$) and ConvSearch (Spatial Distance: $M=5.71$, $SD=1.50$; Temporal Distance: $M=5.47$, $SD=1.60$; Social Distance: $M=6.04$, $SD=1.08$; Hypothetical Distance: $M=5.90$, $SD=1.48$) were not significant (Spatial Distance: $W=52.00$, $p=0.386$, $r=0.33$; Temporal Distance: $t(15)=1.41$, $p=0.262$, Cohen’s $d=0.35$; Social Distance: $t(15)=-2.38$, $p=0.058$, Cohen’s $d=-0.60$; Hypothetical Distance: $W=35.00$, $p=0.366$, $r=-0.33$). The small to moderate effect sizes observed at near distances suggest that the psychological proximity of the task does not strongly influence perceived usefulness between the two search modalities. The positive effect sizes for spatial and temporal distances indicate a slight preference for WebSearch in terms of usefulness, whereas the negative effect sizes for social and hypothetical distances suggest a marginal preference for ConvSearch. Notably, the moderate negative effect size for social distance suggests a trend where participants may begin to perceive ConvSearch as slightly more useful than WebSearch, even at near distances.
@@
In contrast, at a far psychological distance, the preference for ConvSearch showed a significant increase. As depicted in Fig. 5 (e)-(h), substantial differences emerged in favor of ConvSearch across all dimensions. In spatial distance, ConvSearch ($M=6.19$, $SD=0.55$) was perceived as more useful compared to WebSearch ($M=5.04$, $SD=1.48$), $W=2.00$, $p=0.006$, $r=-0.95$. Similar phenomena were observed in the temporal (ConvSearch: $M=6.60$, $SD=0.59$; WebSearch: $M=5.26$, $SD=1.41$; $W=3.00$, $p=0.004$, $r=-0.95$), social (ConvSearch: $M=6.53$, $SD=0.56$; WebSearch: $M=4.85$, $SD=1.57$; $t(15)=-3.66$, $p=0.005$, Cohen’s $d=-0.92$), and hypothetical distances (ConvSearch: $M=6.46$, $SD=0.90$; WebSearch: $M=4.81$, $SD=1.78$; $t(15)=-3.28$, $p=0.008$, Cohen’s $d=-0.82$). The large effect sizes observed at four psychological distances indicate that as the psychological distance of a task increases, participants perceive ConvSearch as significantly more useful than WebSearch. The consistent negative direction of the effect sizes suggests a clear and robust preference for ConvSearch in terms of usefulness under conditions of far psychological distance. Notably, the effect size increased from a moderate level at near distances to a large level at far distances, highlighting the particularly strong influence of psychological distance on perceived usefulness. These findings support H3, demonstrating that as psychological distance increases, so does the perceived usefulness of ConvSearch compared to WebSearch.
5.1.4. Psychological Distance Increases Differences in Perceived Enjoyment (H4 Supported)
@@
Figure 6. Comparative analysis of perceived enjoyment between WebSearch and ConvSearch across four psychological distances. Panels (a)-(c) indicate no significant differences at a near psychological distance, while panel (d) shows a marginal preference for ConvSearch. Panels (e)-(h) demonstrate significant increases in perceived enjoyment for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿ 0.05 (ns), p ¡= 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived enjoyment for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks from the paired samples statistical tests. In the first row, panels (a) through (d) depict the perceived enjoyment at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.514), marked as non-significant (ns); (b) Temporal Distance (p = 0.691), marked as non-significant (ns); (c) Social Distance (p = 0.295), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.041), marked with one asterisk (*). In the second row, panels (e) through (h) illustrate the perceived enjoyment at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.036), marked with one asterisk (*); panels (f) Temporal Distance (p = 0.002), marked with two asterisks (**); (g) Social Distance (p = 0.012), marked with one asterisk (*); (h) Hypothetical Distance (p = 0.006), marked with two asterisks (**).
@@
The analysis of perceived enjoyment revealed significant differences between WebSearch and ConvSearch, particularly as psychological distance increased. At a near psychological distance (as shown in Fig. 6 (a)-(c)), across spatial (WebSearch: $M=5.40$, $SD=1.03$; ConvSearch: $M=5.00$, $SD=1.81$), temporal (WebSearch: $M=5.13$, $SD=1.28$; ConvSearch: $M=4.98$, $SD=1.78$), and social dimensions (WebSearch: $M=5.33$, $SD=1.34$; ConvSearch: $M=5.75$, $SD=1.21$), there were no significant differences in perceived enjoyment (Spatial Distance: $t(15)=0.78$, $p=0.514$, Cohen’s $d=0.19$; Temporal Distance: $t(15)=0.44$, $p=0.691$, Cohen’s $d=0.11$; Social Distance: $t(15)=-1.30$, $p=0.295$, Cohen’s $d=-0.33$). However, a significant difference was noted in the hypothetical distance dimension ($t(15)=-2.65$, $p=0.041$, Cohen’s $d=-0.66$), suggesting a higher perceived enjoyment of ConvSearch ($M=5.67$, $SD=1.22$) compared to that of WebSearch ($M=4.71$, $SD=1.50$) even at a near psychological distance (Fig. 6 (d)). The small effect sizes observed in spatial and temporal distances at near psychological distances suggest only minimal differences in perceived enjoyment, with a slight preference for WebSearch. In contrast, the moderate negative effect size for social distance and the larger negative effect size for hypothetical distance indicate a preference for ConvSearch in terms of enjoyment, particularly in the hypothetical dimension. These findings suggest that even at near distances, certain dimensions of psychological distance (e.g., hypothetical) may influence user preference for ConvSearch.
As psychological distance expanded, differences in perceived enjoyment became more pronounced. In the far psychological distance condition, as depicted in Fig. 6 (e)-(h), all dimensions showed significant differences (Spatial Distance: $t(15)=-2.42$, $p=0.036$, Cohen’s $d=-0.60$; Temporal Distance: $t(15)=-4.73$, $p=0.002$, Cohen’s $d=-1.18$; Social Distance: $t(15)=-3.04$, $p=0.012$, Cohen’s $d=-0.76$; Hypothetical Distance: $W=0.00$, $p=0.006$, $r=-1.00$) favoring ConvSearch (Spatial Distance: $M=5.46$, $SD=1.05$; Temporal Distance: $M=5.92$, $SD=1.23$; Social Distance: $M=5.65$, $SD=1.21$; Hypothetical Distance: $M=6.00$, $SD=1.23$) over WebSearch (Spatial Distance: $M=4.67$, $SD=1.05$; Temporal Distance: $M=4.73$, $SD=1.41$; Social Distance: $M=4.17$, $SD=1.36$; Hypothetical Distance: $M=4.63$, $SD=1.49$). Specifically, the hypothetical distance maintained a higher significant difference than that in near psychological distance. The moderate to large effect sizes observed at far distances for spatial, temporal, social, and hypothetical indicate that as psychological distance increases, ConvSearch is perceived as significantly more enjoyable than WebSearch. The negative direction of the effect sizes consistently reflects a strong preference for ConvSearch in terms of enjoyment. The particularly large effect size for temporal and hypothetical distance suggests that tasks framed with far temporal scenarios (e.g., events occurring in the distant future) or far hypothetical scenarios (e.g., events are unlikely to occur) may elicit the greatest enjoyment from ConvSearch. These results underscore that as psychological distance increases, the perceived enjoyment of using ConvSearch significantly exceeds that of WebSearch across various dimensions, supporting H4.
5.1.5. Psychological Distance Increases Differences in Acceptance of ConvSearch (H5 Supported)
@@
Figure 7. Differential preference for ConvSearch over WebSearch across various psychological distances, shown in two parts: (a) detailed comparison across four dimensions, and (b) aggregated preference at near and far distances. Significance is marked as p ¿ 0.05 (ns), p ¡= 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
@@
The impact of psychological distance on the acceptance of ConvSearch supports H5. When the psychological distance is close (as shown in Fig. 7 (a)), the preference for ConvSearch is already evident, with 75.0% of participants favoring it in the spatial dimension. However, as the distance increases, this preference soars to 87.5%. Similar patterns are observed in other dimensions. For temporal distance, the preference increases from 56.2% at a close distance to 87.5% at a far distance. In social distance, the preference rises from 75.0% to 81.2%, and most notably, in hypothetical distance, the preference jumps from 75.0% to an overwhelming 93.8%. Upon aggregating the data across all dimensions and analyzing the preference for WebSearch and ConvSearch at near and far psychological distances (Fig. 7 (b)), a chi-square test revealed significant differences ( $\chi^{2}(1,N=128)=4.97$, $p=0.026$). This result highlights that while participants exhibit a preference for ConvSearch even at a near psychological distance, this preference intensifies as the psychological distance increases. This indicates that participants showed a significant preference for ConvSearch over WebSearch in scenarios characterized by greater psychological distance.
@@ 5.2.1. Unique Themes for Preferring ConvSearch
The unique thematic of ConvSearch response highlights its distinct advantages over WebSearch.
Low Workload. Many participants (n = 58) reported that using ConvSearch required less effort compared to WebSearch. They felt that ConvSearch simplified the information gathering process by eliminating the need to visit multiple websites and compile information independently. Thus, reduced workload was one of the reasons for their preference for ConvSearch. As one participant explained: “It was easier to find out what I needed to know and less work on my part to pull it all together.” (P6)
Speed and Efficiency. Participants (n = 43) perceived ConvSearch as faster and more efficient in making travel plans. The efficiency they focused on was whether ConvSearch could quickly provide information while meeting their needs to complete travel plans. As one stated, “I would choose the chatbot because it saves me time and is more efficient at making a full-bodied itinerary.” (P108)
Specific Information Consolidation. Another major reason participants (n = 43) preferred ConvSearch was its ability to quickly consolidate information from various sources for specific queries. One participant called it, “The results the chatbot provides is neater compared to the websearch’s results which are just links while the chatbot makes sure to summarize any relevant info and provides it to me.” (P96)